Explain Interest Rate Preferred Habitat Hypothesis Theory and give an example of it?
Answer #1

Yield curve or the term structure of interest rate is a plot of yield to maturity with respect to the time to maturity. There are several theories in finance that attempt to explain the term structure of interest rate. The commonly known theories are;

  1. Segmented markets theory
  2. Expectations theory
  3. Liquidity Preference theory
  4. Preferred habitat theory

Preferred hypothesis theory

Preferred Hypothesis theory makes two assumptions:

  1. Investors are thoughtful of expected return as well as maturity while making investments.
  2. Investors are segmented on the basis of maturity of the instruments. This means that instruments of different maturities will have different sets of investors

If we combine the two assumptions and apply to bond markets or interest rate markets, they can be simply interpreted as investors who are interested in short term maturity bonds will not be willing to hold bonds with longer maturity unless they are paid a premium for holding the investment longer. Short term maturity investors will not prefer long term investments unless they provide incremental return over short term maturity investments.

So, the long term interest rates > short term interest rates. And the difference between the two is called liquidity premium or term premium.

Thus, interest rate on a long term bond = average of short term interest rates occurring over the life of the long term bond + term premium or liquidity premium.

Thus, the investors preferring short term bonds will continue to prefer short term bonds if they are not paid premium for holding long term. This is equivalent to people (short term investors) preferring to stay in their habitat (short term bonds) unless they are incentivised to move out. hence, the theory is called Preferred Habitat Theory.

Let's try to understand the same through an example now.

Suppose one year interest rates over next five years are as follows:

I0,1 = 3%; I1,2 = 4%, I2,3 = 5%; I3,4 = 6% and I4,5 = 7%. Please note that each of these rates are one year interest rate. I1,2 = 4% is to be interpreted as expected interest rate for one year, two years from hence. This also mean one year interest rate, two years from now, is expected to be 4%

Under prefered habitat theory, the interest rate for 2 years maturity will have to be more than the average 1 year interest rate over next two years.

That is annualised I0,2 > (I0,1 + I1,2 ) / 2

That is, I0,2 = (I0,1 + I1,2 ) / 2 + Premium = (3% + 4%) / 2 + Premium = 3.5% + premium

Hence, an investor will be willing to invest in two years maturity only if he / she gets an annualised return of say 3.75% thus reflecting a liquidity premium of 3.75%- 3.5% = 0.25% annually.

Similarly, I0,3 > (I0,1 + I1,2 + I2,3 ) / 3 = (3% + 4% + 5%) / 3 = 4%

So, n investor will be willing to invest in three years maturity only if he / she gets an annualised return of say 4.5% thus reflecting a liquidity premium of 4.5% - 4% = 0.5% annually.